If I have 2(N sq. + 1)-3(2N sq. - 3) and I need to solve for N, how do I work the squared N?

Answer by checkley71(8403)

2(n^2+1)-3(2n^2-3)
2n^2+2-6n^2+9
-4n^2+11 answer.
Unless you set this equation=0. then:
-4n^2+11=0
-4n^2=-11
n^2=-11/-4
n^2=11/4
n=sqrt11/4
n=sqrt2.75
n=1.683 answer.